# Think or Step-by-Step? UnZIPping the Black Box in Zero-Shot Prompts

### v2

@@ Abstract
Zero-shot prompting techniques have significantly improved the performance of Large Language Models (LLMs). However, we lack a clear understanding of why zero-shot prompts are so effective. For example, in the prompt “Let’s think step-by-step,” is think or step-by-step more crucial to its success? Existing interpretability methods, such as gradient-based and attention-based approaches, are computationally intensive and restricted to open-source models. We introduce the ZIP score (Zero-shot Importance of Perturbation score), a versatile metric applicable to both open and closed-source models, based on systematic input word perturbations. Our experiments across four flagship models, seven widely-used prompts, and several tasks reveal that: (1) word importance varies significantly by task type, with mathematical tasks prioritizing structured instruction words; (2) proprietary models align more closely with human judgments than open-source ones; (3) nouns consistently show the highest importance (47.4%–65.9%); and (4) ZIP scores inversely correlate with model performance (|r| > 0.9), indicating stronger prompt effects on challenging tasks. These findings enhance our understanding of LLM behavior and contribute to developing more effective zero-shot prompts and improved model analysis. [^1] The complete codebase and documentation of language model interactions will be publicly available upon publication.
[…]
Think or Step-by-Step?
UnZIPping the Black Box in Zero-Shot Prompts
Nikta Gohari Sadr, Sangmitra Madhusudan, and Ali Emami Brock University, Saint Catharines, Canada {zu22of, sm20pd, aemami}@brocku.ca
@@ 1 Introduction
Prompting techniques have revolutionized how we use Large Language Models (LLMs), significantly improving their performance across various tasks. From Chain-of-Thought (CoT) Wei et al. (2022) to Self-Consistency Wang et al. (2023b) and Least-to-Most prompting Zhou et al. (2022), these methods have expanded LLM capabilities. Zero-shot prompts, which guide models without task-specific training examples, have emerged as particularly effective and practical, offering broad generalization and multi-task abilities Kojima et al. (2022). However, despite their success, we lack a clear understanding of why these prompts are so effective, particularly at the level of individual words.
Consider the following example:
Example 1.
Problem: Alice has 5 apples. She gives 2 to Bob and 1 to Charlie. How many apples does Alice have left?
Zero-shot CoT: Let’s think step-by-step.
Model Output:
  1. 1.
    Alice starts with 5 apples.
  2. 2.
    She gives away 2 + 1 = 3 apples.
  3. 3.
    Apples left = 5 - 3 = 2
Answer: Alice has 2 apples left.
Despite the success of such zero-shot prompting techniques, a gap in our understanding remains: What is it about these prompts that is so effective, particularly at the level of individual words? In the above prompt, which term has more influence on the model output: ‘think’ or ‘step-by-step’?
Current methods for interpreting input word significance in LLMs typically rely on gradients Wallace et al. (2019); Yin and Neubig (2022); Ferrando et al. (2023); Wu et al. (2023) or attention mechanisms Modarressi et al. (2022); Tenney et al. (2020); Ferrando et al. (2022). However, these methods require access to internal model parameters and face reliability issues Wang et al. (2020); Jain and Wallace (2019); Ethayarajh and Jurafsky (2021); Zhao et al. (2024); Bastings and Filippova (2020).
Following established work Choudhary et al. (2022); Saleem et al. (2022); Feng et al. (2018), we define word importance as the degree to which a word’s presence, absence, or modification impacts model performance on a given task. This definition captures both semantic and structural importance through systematic perturbations: a word is considered important if replacing it with semantically similar alternatives (synonyms), related concepts (co-hyponyms), or removing it altogether significantly affects the model’s output.
Our paper introduces the ZIP score (Zero-shot Importance of Perturbation score), a novel metric for assessing word importance in zero-shot instructional prompts that works with both open and closed-source models. Figure 1 provides an overview of the methodology. Our approach is validated through controlled experiments, statistical analyses, and is compared with human judgment across seven zero-shot prompts, four LLMs, and several classification and translation tasks.
@@
Our key findings reveal that: (1) Word importance varies significantly by task type (e.g., "step-by-step" for mathematical tasks vs. "think" for reasoning tasks); (2) Proprietary models like GPT-4o mini align more closely with human judgments of word importance than open-source models; (3) Nouns consistently show the highest importance (47.4%–65.9% of significant words) across all models; and (4) ZIP scores inversely correlate with model performance (|r| > 0.9), indicating that prompts have a greater impact on challenging tasks.
@@ 2 Related Work
Gradient-Based Methods identify influential features by calculating the gradient of the output logits with respect to input elements Wallace et al. (2019). Recent advances include gradient-based post hoc contrastive explanations for model predictions Yin and Neubig (2022); Ferrando et al. (2023) and analyses of how CoT prompting affects saliency scores Wu et al. (2023). However, these methods face significant limitations: they require access to model internals, making them inapplicable to closed-source models. Furthermore, research has shown that model gradients can be easily manipulated, raising concerns about the reliability of gradient-based analyzes Wang et al. (2020).
@@
Perturbation-Based Methods can analyze any LLM regardless of architecture accessibility by measuring how outputs change when inputs are modified. As Choudhary et al. (2022) describes, “a word (token) or a collection of words (tokens) are modified or removed from the input samples, and a resulting change is measured.”
Notable approaches include LIME Ribeiro et al. (2016), which creates local linear approximations of model behavior, and other dataset instance perturbations Zafar and Khan (2019). Recent research has extended these methods to few-shot demonstrations Liu et al. (2023) and system prompts Hackmann et al. (2024); Yin et al. (2023).
@@ 3 The ZIP score
The ZIP score quantifies the importance of individual words in a prompt by measuring how much the model’s performance changes when that word is perturbed or removed. This method allows us to systematically identify which words are most crucial for the model’s performance on a given task.
@@ 3.1 Formalization of the ZIP score
Original Prompt ($P$): “Take a deep breath and work on this problem step-by-step” Luo et al. (2024)
@@
Let $P$ represent the original prompt, where $P=w_{1},w_{2},\dots,w_{n}$ and $w_{i}$ denotes the $i$-th word. [^2] Throughout this paper, we use “word” to refer to any token resulting from our space-based tokenization method. This may include contractions, compound words, or multi-word expressions.
Example:
[…]
Figure 2: Perturbation generation and filtering process for a sample prompt
@@
For each word $w_{i}$ in $P$, we generate a set of perturbations $\{P_{i1},P_{i2},\dots,P_{ik}\}$, through context-aware synonym replacement and co-hyponym replacement using GPT-4333GPT-4 consistently outperformed traditional tools like NLTK Bird et al. (2009) in generating contextually appropriate alternatives in our preliminary tests., and word removal. Synonyms provide similar meanings, co-hyponyms [^4] A co-hyponym is a word that shares the same category or group with another word, but both are specific examples of a broader term (hypernym). introduce related concepts, and removal tests word necessity. Each $P_{ij}$ represents a modified version of $P$ where $w_{i}$ has been replaced or omitted.
[…]
Aligned with existing literature on perturbation-based methods Choudhary et al. (2022); Ivanovs et al. (2021); Saleem et al. (2022), we posit that a word’s importance is indicated by the model’s performance shifts in response to any meaningful alteration of that word, whether through subtle meaningful shifts or its complete removal. (see §7 for analyses and demonstrations of this phenomenon).
To create a final set of perturbations that provides meaningful insight into the impact of each word on the model output, as depicted in Figure 2, we filter the perturbed sentences as follows:
@@
    The Universal Sentence Encoder Cer et al. (2018) to measure semantic similarity between $P$ and $P_{ij}$, keeping those with >30% similarity. [^5] This threshold was chosen to balance maintaining context while allowing meaningful alterations (see table 24 in A.7).
@@
Determined by our filtering process, $P_{42}$ is excluded due to its insufficient semantic similarity and grammatical inconsistencies. Consequently, the final set of perturbations includes $P_{41}$ and $P_{43}$.
We maintained cost-effectiveness in using large language models by optimizing token usage and the output format, reducing the tokens needed for perturbation generation and selection. Additionally, our primary filter, using Universal Sentence Encoder, discards unrelated perturbations, thereby minimizing unnecessary API calls for grammatical verification [^6] For the CoT prompt, the entire perturbation generation and grammatical verification process uses fewer than 2150 input/output tokens, effectively balancing cost and quality.. Across the seven instructional prompts, this filtering process resulted in an average of 9.04 perturbed sentences per word. Detailed prompts and examples for perturbation creation and evaluation are provided in Appendix Sections A.6 and A.7.
@@
We input the original prompt $P$ and each of its validated perturbed variants $P_{ij}$ into the language model $M$ to address an instance of task $T$ which we refer to as $t$.
@@
[…]
@@
We define a disagreement function $d$ that measures the difference between the model predictions for each perturbed prompt ($P_{ij}$) and the original prompt ($P$). This function varies according to the type of task $T$. For instance $t$ of classification tasks:
@@
Example:
[…]
For an instance $t$ of translation tasks:
@@
Translation Example: For the task of translating “The cat is on the mat” to French, assume:
$\text{BLEU}_{t}^{M}(P)=0.85$,
$\text{BLEU}_{t}^{M}(P_{41})=0.83$,
$\text{BLEU}_{t}^{M}(P_{43})=0.72$
$d_{t}^{M}(P,P_{41})=|0.85-0.83|=0.02$, $d_{t}^{M}(P,P_{43})=|0.85-0.72|=0.13$
Step 5: ZIP score Calculation Finally, for each word $w_{i}$, we compute the ZIP score $\text{ZIP}_{T}^{M}(w_{i})$ by averaging disagreement scores across all perturbations of $w_{i}$ and all dataset instances:
$\text{ZIP}_{T}^{M}(w_{i})=100\cdot\frac{1}{N}\sum_{t=1}^{N}\left(\frac{1}{k}\sum_{j=1}^{k}d_{t}^{M}(P_{i},P_{ij})\right)$
Where $N$ is the number of dataset instances, $k$ is the number of valid perturbations for each word, $P_{i}$ is the original prompt for the $i$-th dataset instance, and $P_{ij}$ is the $j$-th perturbation of word $w_{i}$ in the $t$-th dataset instance.
[…]
$\text{ZIP}_{Classif.}^{M}(\text{``breath''})=100\cdot\frac{1}{2}(0+1)=50$
$\text{ZIP}_{Transl.}^{M}(\text{``breath''})=100\cdot\frac{1}{2}(0.02+0.13)=7.5$
The ZIP score ($\in[0,100]$) [^7] The 100x scaling factor normalizes ZIP scores to a 0-100 range for easier interpretation and comparison across tasks. represents the average percentage change in model performance observed across all tested dataset instances when perturbing a given word. [^8] For simplicity, we show the process for one word in a dataset instance; in practice, this repeats for all prompt words, averaging ZIP scores across instances. This metric is adaptable to various performance measures, making it versatile for different types of tasks. Interpreting the ZIP score is straightforward: a higher score indicates greater word importance, as perturbing it leads to larger changes in model output. Conversely, a lower score suggests the word has less impact on performance for the given task.
@@ 3.2 Identifying Significantly Important Words
    For each word in the prompt, we generate a set of perturbed prompts (the same set used in the ZIP score calculation).
@@
    We repeatedly input both the original prompt and each of its perturbed versions to the model, generating multiple outputs for each. The number of times we re-prompt with the original prompt matches the number of perturbations for that word, ensuring a balanced comparison.
@@
The Wilcoxon rank-sum test evaluates whether two independent samples (i.e., the outputs from the original re-prompts and a perturbed prompt) come from the same distribution. Words yielding a p-value < 0.05 are considered ‘significantly important’.
4 Controlled Pre-Experiment

### v3

# Which Words Matter Most in Zero-Shot Prompts?

@@ Abstract
While zero-shot instructional prompts like “Let’s think step-by-step” have revolutionized Large Language Model performance, a fundamental question remains unanswered: which specific words drive their remarkable effectiveness? We introduce the ZIP score (Zero-shot Importance of Perturbation), the first systematic method to quantify individual word importance in instructional prompts through controlled perturbations including synonym replacement, co-hyponym substitution, and strategic removal. Our analysis across four flagship models, seven widely-adopted prompts, and multiple task domains reveals four key findings: (1) Task-specific word hierarchies exist where mathematical problems prioritize “step-by-step” while reasoning tasks favor “think”; (2) Proprietary models show superior alignment with human intuitions compared to open-source alternatives; (3) Nouns dominate importance rankings, consistently representing the majority of significant words; and (4) Word importance inversely correlates with model performance, indicating prompts have greatest impact where models struggle most. Beyond revealing these patterns, we establish the first ground-truth benchmark for prompt interpretability through 20 validation prompts with predetermined key words, where ZIP achieves 90% accuracy versus LIME’s 60%. Our findings advance prompt science, the study of how language shapes model behavior, providing both practical insights for prompt engineering and theoretical understanding of word-level effects in LLMs. [^1] The complete codebase and documentation of language model interactions will be publicly available upon publication.
Which Words Matter Most in Zero-Shot Prompts?
Nikta Gohari Sadr1, Sangmitra Madhusudan1, Hassan Sajjad2 and Ali Emami3 1Brock University, St. Catharines, Canada 2Dalhousie University, Halifax, Canada 3Emory University, Atlanta, USA
@@ 1 Introduction
Chain-of-Thought prompting transformed how we use Large Language Models. Adding “Let’s think step-by-step” to a prompt, GPT-3 suddenly solved complex math problems it previously failed on Wei et al. (2022). This discovery sparked a wave of instruction prompts: Self-Consistency Wang et al. (2023b), Plan-and-Solve Wang et al. (2023a), and others, each improving performance through carefully crafted phrases. These prompts didn’t just boost task performance; they became fundamental to training Large Reasoning Models like OpenAI’s o3 OpenAI (2025).
Consider the Plan-and-Solve prompt: “Let’s first understand the problem and devise a plan to solve it. Then, solve it step-by-step” Wang et al. (2023a). This manually crafted extension of Chain-of-Thought consistently outperforms the original prompt across benchmarks. The intuition seems clear: explicit planning should reduce errors. But which words actually drive this improvement? Is it the instruction to “plan”? The directive to “understand”? Or does the familiar “step-by-step” do most of the work? Despite the empirical success of these prompts, the selection of which words to include remains largely ad hoc, we lack systematic understanding of how individual word choices affect model performance.
While it is well-known that models are sensitive to prompt structure Sclar et al. (2024), the importance of individual word choices in a prompt remains poorly understood. Individual word choices can have unexpected and counterintuitive effects on model performance. For instance, replacing “solve” with its synonym “work out” seems harmless, but can produce vastly different outputs.
Following established work on perturbation-based interpretability Choudhary et al. (2022); Saleem et al. (2022); Feng et al. (2018), we define word importance as the degree to which a word’s presence, absence, or modification impacts model performance on a given task. While gradient-based Wallace et al. (2019); Yin and Neubig (2022); Ferrando et al. (2023); Wu et al. (2023) and attention-based methods Modarressi et al. (2022); Tenney et al. (2020); Ferrando et al. (2022) also aim to identify important words, they infer importance from model internals rather than directly measuring performance impact.
Beyond this, methods that analyze internal representations face practical limitations. They require access to model internals, making them inapplicable to closed-source models like GPT-4. They also face reliability issues: gradients can be manipulated Wang et al. (2020) and attention weights often fail to correlate with actual performance impact Jain and Wallace (2019); Ethayarajh and Jurafsky (2021). Recent work on mechanistic interpretability (i.e., circuit analysis, sparse autoencoders, and activation patching) Lindsey et al. (2025) provides detailed internal analysis but requires substantial computational resources and whitebox access. For understanding how word choices actually affect model outputs, we need lightweight, model-agnostic methods that directly measure performance changes.
We introduce the ZIP score (Zero-shot Importance of Perturbation score), a metric that quantifies word importance through systematic perturbations including synonym replacement, co-hyponym substitution, and word removal. Our approach works with any model, requiring only the ability to query the model. Additionally, we develop a set of validation prompts with known ground-truth word importance, providing a benchmark for evaluating interpretability methods. Figure 1 shows our method identifying “step-by-step” as the critical component in Chain-of-Thought prompting for a mathematical reasoning task.
@@
Testing across four flagship models, seven prompts, and multiple tasks reveals surprising patterns. Word importance varies dramatically by task type: “step-by-step” dominates in mathematical problems while “think” matters more for common-sense reasoning. Proprietary models like GPT-4o mini show closer alignment with human intuitions than open-source alternatives. Across all models, nouns consistently emerge as most important (47.4%–65.9% of significant words). Most strikingly, ZIP scores inversely correlate with model performance $(|\text{r}|>0.9)$: prompts have their greatest impact on tasks where models struggle.
These findings advance prompt science Shah (2025), the systematic study of how prompts elicit model behavior, by quantifying word level importance in zero-shot instructional prompts. By revealing which words actually matter and how word choices affect model outputs, we provide both practical guidance for prompt design and deeper insights into prompt-model interactions.
@@ 2 Related Work
Gradient-Based Methods identify influential features by calculating the gradient of the output logits with respect to input elements Wallace et al. (2019). Recent advances include gradient-based post hoc contrastive explanations for model predictions Yin and Neubig (2022); Ferrando et al. (2023) and analyses of how CoT prompting affects saliency scores Wu et al. (2023). However, these methods face significant limitations: they require access to model internals, making them inapplicable to closed-source models. Furthermore, research has shown that model gradients can be easily manipulated, raising concerns about the reliability of gradient-based analyses Wang et al. (2020).
@@
Perturbation-Based Methods can analyze any LLM regardless of architecture accessibility by measuring how outputs change when inputs are modified. As Choudhary et al. (2022) describes, “a word (token) or a collection of words (tokens) are modified or removed from the input samples, and a resulting change is measured.” Notable approaches include LIME Ribeiro et al. (2016), which creates local linear approximations of model behavior, and other dataset instance perturbations Zafar and Khan (2019). Recent research has extended these methods to few-shot demonstrations Liu et al. (2023) and system prompts Hackmann et al. (2024); Yin et al. (2023).
@@
Prompt Science: Recent work distinguishes between prompt engineering (optimizing prompts for specific tasks) and prompt science (using prompts to discover regularities in model behavior) Holtzman and Tan (2025); Shah (2025). Studies have documented prompt brittleness, the sensitivity of model outputs to minor prompt variations Arora et al. (2022); Min et al. (2022); Sclar et al. (2023). For instance, Sclar et al. (2023) found that formatting changes alone can swing accuracy by over 70 percentage points. Rather than viewing this sensitivity as a flaw, we leverage it as a tool for understanding word importance. As Holtzman and Tan (2025) argued, prompt sensitivity reflects models’ attempts to “infer substantial information from limited context,” making it a feature for scientific investigation rather than a bug to be fixed. Our ZIP score provides a systematic framework for what they call “behavioral studies that use varied prompts in structured ways to confirm hypotheses.”
@@ 3 The ZIP score
The ZIP score quantifies the importance of individual words in a prompt by measuring how much the model’s performance changes when that word is perturbed or removed.
@@ 3.1 Formalization of the ZIP score
Original Prompt ($P$): “Take a deep breath and work on this problem step-by-step” Luo et al. (2024).
@@
Let $P$ represent the original prompt, where $P=w_{1},w_{2},\dots,w_{I}$ and $w_{i}$ denotes the $i$-th word. Throughout this paper, we use “word” to refer to any token resulting from our space-based tokenization method. This may include contractions, compound words, or multi-word expressions. Example:
$P=[\text{``Take''},\text{``a''},\text{``deep''},\ldots,\text{``step-by-step''}]$
Figure 2: Perturbation generation and filtering pipeline
@@
For each word $w_{i}$ in $P$, we generate a set of perturbations $\{P_{i1},P_{i2},\dots,P_{iJ}\}$, through context-aware synonym replacement and co-hyponym replacement using GPT-4. Synonyms provide similar meanings, co-hyponyms introduce related concepts, and removal tests word necessity. Each $P_{ij}$ represents a modified version of $P$ where $w_{i}$ has been replaced or omitted.
Example: For $w_{4}=\text{``breath''}$:
  • $P_{41}=\text{``Take a deep pause and work on ...''}$
  • $P_{42}=\text{``Take a deep glimpse and work on ...''}$
  • $P_{43}=\text{``Take a deep look and work on ...''}$
Aligned with existing literature on perturbation-based methods Choudhary et al. (2022); Ivanovs et al. (2021); Saleem et al. (2022), we posit that a word’s importance is indicated by the model’s performance shifts in response to any meaningful alteration of that word, whether through subtle meaningful shifts or its complete removal (see §6 for analyses and demonstrations).
To create a final set of perturbations that provides insight into the impact of each word on the model output, as depicted in Figure 2, we filter the perturbed sentences as follows:
@@
    The Universal Sentence Encoder Cer et al. (2018) to measure semantic similarity between $P$ and $P_{ij}$, keeping those with >>30% similarity. This threshold was chosen to balance maintaining context while allowing meaningful alterations (see Table 21 in A.7).
@@
Our filtering process excludes $P_{42}$ for insufficient semantic similarity and grammatical issues, leaving $P_{41}$ and $P_{43}$ as valid perturbations.
We maintained cost-effectiveness by designing concise prompts and compact output formats to minimize token usage, and by using the Universal Sentence Encoder to filter out unrelated perturbations before making API calls. Across the seven instructional prompts, this filtering process resulted in an average of 9.04 perturbed sentences per word. Detailed prompts and examples for perturbation creation and evaluation are provided in Appendix Sections A.6 and A.7.
@@
We input the original prompt $P$ and each of its validated perturbed variants $P_{ij}$ into the language model $M$ to address an instance $(t)$ of task $T$.
@@
$\text{pred}_{t}^{M}(P)$ $=\text{``Prime''}\quad\text{(Correct)}$
$\text{pred}_{t}^{M}(P_{41})$ $=\text{``Prime''}\quad\text{(Correct)}$
$\text{pred}_{t}^{M}(P_{43})$ $=\text{``Composite''}\quad\text{(Incorrect)}$
@@
We define a disagreement function $d$ that measures the difference between model predictions for each perturbed prompt ($P_{ij}$) and the original prompt ($P$). This function varies by task type $T$. In our study, for classification tasks:
@@
For translation tasks:
@@
Since the above example $T$ is a classification task, we calculate disagreement scores as follows:
$d_{t}^{M}(P,P_{41})$ $=0\quad\text{(No disagreement)}$
$d_{t}^{M}(P,P_{43})$ $=1\quad\text{(Disagreement)}$
Step 5: ZIP score Calculation
Finally, for each word $w_{i}$, we compute $\text{ZIP}_{T}^{M}(w_{i})$ by averaging disagreement scores across all $J$ perturbations of $w_{i}$ and all dataset instances:
$\text{ZIP}_{T}^{M}(w_{i})=100\cdot\frac{1}{N}\sum_{t=1}^{N}\left(\frac{1}{J}\sum_{j=1}^{J}d_{t}^{M}(P,P_{ij})\right),$
where $N$ is the number of dataset instances, $P$ is the original prompt, and $P_{ij}$ is the $j$-th perturbation of word $w_{i}$ in the $t$-th dataset instance.
Example: For $w_{4}=\text{``breath''}$ in a single instance ($N=1$):
$\text{ZIP}_{T}^{M}(\text{``breath''})=100\cdot\frac{1}{2}(0+1)=50$
The ZIP score ($\in[0,100]$) represents the average percentage change in model performance observed across all tested dataset instances when perturbing a given word. This metric is adaptable to various performance measures, making it versatile for different types of tasks. Interpreting the ZIP score is straightforward: a higher score indicates greater word importance, as perturbing it leads to larger changes in model output. Conversely, a lower score suggests the word has less impact on performance for the given task.
Code Prompt Mutual Datasets
0-CoT Let’s think step-by-step. Kojima et al. (2024) BIG-bench, GSM8K, AQUA-RAT
0-CoTB Take a deep breath and work on this problem step-by-step. Yang et al. (2024) AQUA-RAT, GSM8K
0-CoTR Let’s work this out in a step-by-step way to be sure we have the right answer. Zhou et al. (2023) GSM8K, BIG-bench
0-IRR Feel free to ignore irrelevant information in the problem description. Shi et al. (2023) GSM8k
0-PS Let’s first understand the problem and devise a plan to solve it. Then, solve it step-by-step. (Plan & Solve Wang et al. (2023a)) GSM8K, AQUA-RAT
0-DSP Provide the translation step-by-step, then complete the sentence. Peng et al. (2023) WMT19 (German, Chinese)
0-DTG Detect the error type first, then refine the translation. (Deliberate then Generate Li et al. (2023)) WMT (German, Chinese)
Table 1: Zero-shot instructional prompts used in our experiments, and mutual datasets they were tested on.
@@ 3.2 Identifying Significantly Important Words
    For prompt word, we generate a set of perturbed prompts (the same set used in the ZIP score calculation).
@@
    We repeatedly input both the original prompt and each of its perturbed versions to the model, generating multiple outputs for each, up to the maximum number of word perturbations in the prompt.
@@
The Wilcoxon rank-sum test tests whether two independent samples (i.e., the outputs from the original re-prompts and a perturbed prompt) come from the same distribution. We deem words with a p-value << 0.05 as ‘significantly important’.
4 Experimental Setup
4.1 Validation Prompts